OpenStudy (anonymous):

use substitution to evaluate the limit limit [(1/h)/( ln (1+1/h))] as h goes to infinity

OpenStudy (anonymous):

\[\lim_{h \rightarrow \infty} \frac{\frac{1}{h}}{1 + \frac{1}{h}}\]

OpenStudy (anonymous):

what

OpenStudy (anonymous):

Just formatting the question to make sure it's the one you're asking

OpenStudy (anonymous):

its 1/h / ln1+ 1/h you missed the ln (natural log)

OpenStudy (anonymous):

L'Hopital

OpenStudy (anonymous):

havent learned lhosp..

OpenStudy (turingtest):

\[\lim_{h \rightarrow \infty} \frac{\frac{1}{h}}{\ln(1 + \frac{1}{h})}\]right?

OpenStudy (turingtest):

\[\lim_{h \rightarrow \infty} \frac{\frac{1}{h}}{\ln(1 + \frac{1}{h})}\]let \(t=\frac1h\) then as \(h\to\infty\) we have that \(t\to0\) so we get\[\lim_{h \rightarrow \infty} \frac{\frac{1}{h}}{\ln(1 + \frac{1}{h})}=\lim_{t\to0}\frac t{\ln(1+t)}\]hm... how to do this without l'hospital...

OpenStudy (anonymous):

thats my problem lol

OpenStudy (anonymous):

wasnt sure if to problem the h from the 1/h down..

OpenStudy (anonymous):